OpenStudy (anonymous):

Find the midpoint between two points on a coordinate plane whose coordinates are (5, 7) and (-3, 1).

OpenStudy (tkhunny):

Have you considered the simple average of the given coordinates?

OpenStudy (anonymous):

2, 8) (1, 4) (8, 8) (4, 4)

OpenStudy (tkhunny):

Yeah, I almost never care what the answers are. You didn't answer my question.

OpenStudy (mathmale):

Another way in which you might approach this problem: First, take the two x-coordinates and graph them on the x-axis. Find the coordinate of the point that is exactly halfway between the two x-coordinates. That coordinate will the be x-coordinate of the midpoint. Were you to follow tkhunny's suggestion and find the average of these two x-coordinates, you'd necessarily have to come up with the same result. Your choice. Now find the y-coordinate of the midpoint, in the same manner.

OpenStudy (tkhunny):

Example. Given 2 and 4 ==> \(\dfrac{2+4}{2} = \dfrac{6}{2} = 3\) - This is the "Simple Average".

OpenStudy (anonymous):

@tkhunny you and your little smart comments can get tf on honestly!

OpenStudy (tkhunny):

Innocent on this on. I wondered if that name for that process would be a problem. Really, it's just called a "Simple Average". There are other kinds of averages that are more complicated. That's just it's name. I am not, in any way, saying it is "simple" and you should feel bad about not getting it on the first try.